OpenStudy (calculusxy):

A rug is to be completely trimmed with fringe. If the rug is 32 square inches and its length is twice its width, how many inches of fringe will be needed for the rug? A) 18 B) 24 C) 30 D) 36 E) 42

imqwerty (imqwerty):

u done with this ques?

OpenStudy (calculusxy):

No it accidentally or just suddenly closed.

imqwerty (imqwerty):

haha ok :D

imqwerty (imqwerty):

let the width of rug be w its length will be=2w its area=length x width 32=w(2w) can u find w and then the length (2w)

OpenStudy (calculusxy):

2w^2 = 32 w^2 = 16 w = 4 And 2w = 2(4) = 8

imqwerty (imqwerty):

yes so now u have the length and width u only need to find the length of fringe so thas the border length=perimeter of rectangular rug so can u find this :)